Solve the following boundary value problem
Ut=Uxx,0<x<l,t>0
U(0,t)=U(l,t)=0
U(x,0)=x(l-x),0<=x<=1
1
Expert's answer
2019-04-22T10:22:20-0400

Separate variables. Let

"U(x, t)=X(x)T(t)"

Then


"U_t=XT', U_{xx}=X''T"

Substitute


"XT'=X''T"

Dividing this equation by "XT", we have


"\frac{T'}{T}=\frac{X''}{X}=-\lambda"

for some constant "\lambda". Therefore, if there exists a solution "U(x, t)=X(x)T(t)" of the heat equation, then "X(x)" and "T(t)" must satisfy the equations 

"\frac{T'}{T}=-\lambda"

"\frac{X''}{X}=-\lambda"

Solve the eigenvalue problem


"X''+\lambda X=0, X(0)=X(1)=0"

We have to find nontrivial solutions X of the eigenvalue problem. If λ=0, then "X(x)=\alpha x+\beta" and "0=X(0)=0+\beta" implies that "\beta=0" and "0=X(1)=\alpha" implies that "\alpha=0". If "\lambda<0" then "0=X(0)=\alpha\cosh(0)+\beta\sinh(0)=\alpha" and "0=X(1)=0+\beta\sinh(1)=\beta\sinh(1)" shows that also "\beta=0". We conclude that "\lambda\leq0"  is not an eigenvalue of the problem. Finally, consider "\lambda>0." Then "0=X(0)=\alpha\cos(\sqrt{\lambda}\cdot0)+0=\alpha" and "0=X(1)=\beta\sin(\sqrt{\lambda}\cdot1)."

Since "X" is nontrivial solution, "\beta=\not0" and hence "\sin(\sqrt{\lambda}\cdot1)=0." Consequently,


"\lambda=(\frac{n\pi}{1})^2, n\geq1\pi\pi"

and the corresponding eigenfunction is given by


"X_n(x)=\sin(\frac{n\pi x}{1})"

After substituting "\lambda=(\dfrac{n\pi }{1})^2" , we get the family of solutions


"T_n(t)=B_ne^{-k({n\pi \over 1})^2t}"

Thus we have obtained the following sequence of solutions


"U_n(x, t)=X_n(x)T_n(t)=B_n\sin(\frac{n\pi x}{1})e^{-k({n\pi \over 1})^2t}"

We obtain more solutions by taking linear combinations of the "U_n" ’s ( recall the superposition principle) 


"U(x, t)=\displaystyle\sum_{n=1}^NX_n(x)T_n(t)=\displaystyle\sum_{n=1}^NB_n\sin(\frac{n\pi x}{1})e^{-k({n\pi \over 1})^2t}"

and then by passing to the limit "N\to \infin,"


"U(x, t)=\displaystyle\sum_{n=1}^{\infin}X_n(x)T_n(t)=\displaystyle\sum_{n=1}^{\infin}B_n\sin(\frac{n\pi x}{1})e^{-k({n\pi \over 1})^2t}"

Finally, we consider the initial condition. At "t=0", we must have


"U(x, 0)=\displaystyle\sum_{n=1}^{\infin}B_n\sin(\frac{n\pi x}{1})=x(1-x)"

"B_n={2 \over 1}\displaystyle\int_{0}^1x(1-x)\sin(\frac{n\pi x}{1})dx"

"\int xsin(n\pi x)dx=-{1 \over \pi n}xcos(n\pi x)+{1 \over \pi n}\int cos(n\pi x)dx="

"=-{1 \over \pi n}xcos(n\pi x)+{1 \over \pi^2 n^2}sin(n\pi x)+C_1"

"-\int x^2sin(n\pi x)dx={1 \over \pi n}x^2cos(n\pi x)-2{1 \over \pi n}\int xcos(n\pi x)dx"

"-2{1 \over \pi n}\int xcos(n\pi x)dx=-2{1 \over \pi^2 n^2}xsin(n\pi x)+2{1 \over \pi^2 n^2}\int sin(n\pi x)dx="

"=-2{1 \over \pi^2 n^2}xsin(n\pi x)-2{1 \over \pi^3 n^3}cos(n\pi x)+C_2"

"B_n=2[-{1 \over \pi n}xcos(n\pi x)+{1 \over \pi^2 n^2}sin(n\pi x)]\begin{matrix}
 1 \\
 0
\end{matrix}+"

"+2[{1 \over \pi n}x^2cos(n\pi x)-{2 \over \pi^2 n^2}xsin(n\pi x)-{2 \over \pi^3 n^3}cos(n\pi x)]\begin{matrix}
 1 \\
 0
\end{matrix}="

"=-{2 \over \pi n}cos(n\pi )+{2 \over \pi n}cos(n\pi )-{4 \over \pi^3 n^3}cos(n\pi )+{4 \over \pi^3 n^3}="


"={4(1-(-1)^n) \over \pi^3 n^3}"

We arrive at the solution


"U(x, t)=\displaystyle\sum_{n=1}^{\infin}X_n(x)T_n(t)=\displaystyle\sum_{n=1}^{\infin}{4(1-(-1)^n) \over \pi^3 n^3}\sin(\frac{n\pi x}{1})e^{-k({n\pi \over 1})^2t}"